Home‎ > ‎Crew 165‎ > ‎

0308-DailySummaryReport

0308-DailySummaryReport

Daily Summary Report (DSR) for 08 March 2016 and our Sol 2 on Mars!
SOL: 2
Person filling out Report: Ken Sullivan
Summary Title: NorCal Rover Testing, photo video
Mission Status: All systems nominal

Sol 2 Activity Summary: Food measurements are becoming more routine, but weighing ourselves before and after going to the bathroom seems odd and out of place. Body is ready to poo and I have to wait another minute to weigh myself and do paperwork.  Hope diarrhea doesn't
show up. 

Anomalies in work: None
 
Weather: Mostly sunny, light wind.  

Look ahead plan: Research, cleaning and catching up as there will be no EVA
tomorrow. 

Crew Physical Status: Good!

EVA: EVA 2 success. More garbage collection and good photo's and video's.  NorCal rover initial trials were good although signal was lost and ended the rover operations. 

Reports to be filed:
- Commander Report
- Engineering Report
- Journalist Report
- Daily Summary Report
- EVA #2 Report

Support Requested:
Water!  Zero water in trailer tank and static tank is now 13.5  inches from the bottom (which doesn't include unusable water).  We have no ATV fuel in the two 5 gallon containers.